Pa got in his '27 ford and started driving west from Memphis at a constant speed of 40 mph. An hour later ma got in her '34 Buick and started driving west from Memphis at a constant speed. After driving 4 hours ma had passed pa and was 20 miles ahead of him. How fast was ma driving?

Answer by ankor@dixie-net.com(22740)

Pa got in his '27 ford and started driving west from Memphis at a constant speed of 40 mph.
An hour later ma got in her '34 Buick and started driving west from Memphis at a constant speed.
After driving 4 hours ma had passed pa and was 20 miles ahead of him. How fast was ma driving?
:
From the information given, Pa's travel time = 5 hrs, Ma's = 4 hrs
Let s = Ma's driving speed
:
4s - 20 = 5(40)
4s = 200 + 20
s = 220/4
s = 55 mph is Ma's speed
:
:
See if that checks out, find the distance each traveled
4 * 55 = 220
5 * 40 = 200
------------
differ: 20 mi